OpenStudy (anonymous):

Order 34 x 10^2, 1.2 x 10^-7, 8.11 x 10-3, and 435 from least to greatest.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 435 = 4.35 * 10^2 }\) The least power is the least and the greatest is greatest.

OpenStudy (unklerhaukus):

\[34 \times 10^2=3400,\]\[ 1.2 \times 10^{-7}=0.00000012,\]\[8.11 \times 10^{-3}=0.00811, \]\[\qquad\qquad435=435\] can you order them now?